# An axially symmetric stationary N-center solution of Einstein’s vacuum equations

### 09 Mar 2026 v1

@@ 5 Two-stationary euclidon (soliton) solution (the Kerr-NUT solution)
b) It turns out possible to generalize the two-stationary euclidon (soliton) solution 5.3 with an arbitrary stationary vacuum field $\varepsilon^{0}=f^{0}+i\Phi^{0}$ in the Weyl canonical coordinates $(\rho,z)$. For this purpose it is better to use a single relation for a complex function $\epsilon=f+i\Phi$, i.e. [2]
@@
and the functions $\ln a=-U^{a},\ \ \ln b=U^{b}$ should be found from the first-order differential equations 4.5 and $U^{a}=U_{z_{2},f^{0}}(\rho,z),\ \ U^{b}=U_{z_{1},f^{0}}(\rho,z)$.
@@ 6 N-center solution
$U=U_{1:z_{1};n}(\rho,z)=\sum^{2n}_{i=1}(-1)^{i+1}\gamma_{i}\ln\frac{R_{i}+r_{1}+z_{1}-Z_{i}}{R_{i}-r_{1}+z_{1}-Z_{i}},$ (6.3)
@@
$U=U_{1:z_{1};n}(\rho,z)=\sum^{2n}_{i=1}(-1)^{i+1}\gamma_{i}\ln\frac{r_{i}+r_{1}+z_{1}-z_{i}}{r_{i}-r_{1}+z_{1}-z_{i}}=$
$\sum^{2n}_{i=1}(-1)^{i+1}\gamma_{i}\ln\frac{(r_{i}+r_{1}+z_{1}-z_{i})^{2}}{2(z_{1}-z_{i})(z-z_{i}+r_{i})},$ (6.14)
@@ 7 Two-center solution
$f_{2:k_{0},-k_{0};1}\to 1-\frac{2\delta m}{r},\ \ \omega_{2:k_{0},-k_{0};1}\to\frac{\delta J\sin^{2}{\theta}}{r},$ (7.6)
@@
Tomimatsu-Sato solution [136] for deformation parametr $\delta=2$ has has the same asymptotics 7.6.

### 13 Mar 2026 v2

@@ 5 Two-stationary euclidon (soliton) solution (the Kerr-NUT solution)
b) It turns out possible to generalize the two-stationary euclidon (soliton) solution 5.3 with an arbitrary stationary vacuum field $\varepsilon^{0}=f^{0}+i\Phi^{0}$ in the Weyl canonical coordinates $(\rho,z)$. For this purpose it is better to use a single relation for a complex function $\varepsilon=f+i\Phi$, i.e. [2]
@@
and the functions $a=\exp{(-U^{a})},\ \ b=-\exp{U^{b}}$ should be found from the first-order differential equations 4.5 and $U^{a}=U_{z_{2},f^{0}}(\rho,z),\ \ U^{b}=U_{z_{1},f^{0}}(\rho,z)$.
@@ 6 N-center solution
$U=U_{1:z_{1};n}(\rho,z)=\sum^{2n}_{i=1}(-1)^{i+1}\gamma_{i}\ln\frac{R_{i}+r_{1}+z_{1}-Z_{i}}{R_{i}-r_{1}+z_{1}-Z_{i}}+C,$ (6.3)
$C$ is an integrating constant.
@@
$U=U_{1:z_{1};n}(\rho,z)=\sum^{2n}_{i=1}(-1)^{i+1}\gamma_{i}\ln\frac{r_{i}+r_{1}+z_{1}-z_{i}}{r_{i}-r_{1}+z_{1}-z_{i}}+C=$
$\sum^{2n}_{i=1}(-1)^{i+1}\gamma_{i}\ln\frac{(r_{i}+r_{1}+z_{1}-z_{i})^{2}}{2(z_{1}-z_{i})(z-z_{i}+r_{i})}+C,$ (6.14)
$C$ is an integrating constant.
@@ 7 Two-center solution
$f_{2:k_{0},-k_{0};1}\to 1-\frac{2\delta m}{r},\ \ \omega_{2:k_{0},-k_{0};1}\to\frac{2\delta J\sin^{2}{\theta}}{r},$ (7.6)
@@
Tomimatsu-Sato solution [136] for deformation parametr $\delta=2$ has the same asymptotics 7.6.